ITED + atalog + + fers EXPENSES (Thousands of dollars per day) 0 r Network 0 10 20 30 70 80 90 100 40 50 60 TIME (Days)

Math and Graphing Assessment with Tutorials (Mobile-Enabled) 

The area of the shaded quadrilateral is _______

The area of the quadrilateral corresponds to the _______ between _______
Answer #1

Rezercet) (Thousands of $ ----- - - Expenser 27 - - - 20 Time (Dam) Area of Quadrilateral = area Ae area B Area A = = area ofpиe и od 2 -c (so, 6) - - - p (20, 2) А (20,о) В (Go, o) area of quonde lateral ABCD = area of ABD + area of дверArea of by the ABD and concept of ABCD can be determinent found { f(x, , 92 s( X2, 9,) 8 (3, 9а) оча three non- wllinean poinалое оf 4 В СФ В(Goo) C ( 60, 62, p(202) 5 | 26 2 6o (6-2) – о + ( 120 = ( 24o +о) – туҷо = 12-о area of quadrilateral ABCD =